What mass of Na2S2O3 is needed to react with 0.12 g of Cl2?
Step 2: Mass to moles
Step 3: Mole to Mole (Ratio)
Step 4: Mole to Mass
1
Expert's answer
2020-12-06T10:34:35-0500

Na2S2O3 + Cl2 + H2O = S + Na2SO4 + 2 HCl

M (Cl2) = 70.91 g/mol

M (Na2S2O3) = 158.11 g/mol

n (Cl2) = m/M = 0.12 / 70.91 = 0.002 mol

n (Na2S2O3) = n (Cl2) = 0.002 mol

m (Na2S2O3) = n x M = 0.002 x 158.11 = 0.32 g